I seem to be having trouble with a few questions my main one being.....two cars leave Denver at the ame time and travel in opposite directions. One car travels 10mi/h faster than the other car. The cars are 500mi apart in 5 hr. How fast is each car traveling? i would really appreciate some help i dont want just the answer i want help with it and for it to be explained? PLEASE!!!!
Answer by checkley77(12844)

D=RT
500=(X+X+10)5
500=(2X+10)5
500=10X+50
10X=500-50
10X=450
X=450/10
X=45 MPH FOR THE SLOWER CAR.
45+10=55 MPH FOR THE FASTER CAR.
PROOF:
500=(45+55)5
500=100*5
500=500